OpenStudy (anonymous):

two car are 250 miles apart and are traveling towards each other. the first car is traveling 15 mph slower than the second car. after two hours the cars pass each other. how fast is each car traveling?

OpenStudy (anonymous):

Oa------>(x-15) (x)<--------Ob |---------------250-----------------| With the equation S=So+Vo(t), we can find: Sa = Soa + Voa(t) Sb = Sob + Voa(t) Sa = 0 + (x-15)(t) Sb = 250 - x(t) Sa = (x-15)(t) after 2hr: Sa = (x-15)(2) Sb = 250 - x(2) Sa = 2x-30 Sb = 250 - 2x When the cars pass each other, it means that Sa = Sb. Therefore, 2x-30 = 250-2x 4x=280 x=70 Car a: Car b: x-15 => 70-15 = 55 x = 70 is traveling at 55mph is traveling at 70mph Is that right?

OpenStudy (anonymous):

yea but is there an easier way to solve it

OpenStudy (anonymous):

not that i know of